• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Al Tobias

    Log entry time 08:11:18 on December 17, 2009

    Entry number 307624

    keyword=Owl Shift Summary

    
    
    Owl Shift Summary:  A. Tobias (SL)  K. Meyers (TO)  J. Rubin (3rd)
    ==================================================================
    
    EVENTS
    ------
    
    00:00  Shift starts with continued production data on LD2 Slug 24.
    
    01:25  Compton DAQ ROCs not responding after I initial an EndRun.
           Coda reboot, Download transition failed. Coda reboot 2nd
           time, start run 21636 at 01:36 OK.
    
    01:26  Beam trips off and stay off for about half an hour.
    
    01:55  Beam is back.
    
    04:30  Kai finished trying to diagnose and solve an issue with
           connection to TS1 and ROC23 of the PVDIS DAQ. Seems cannot
           telnet into hatsv9. Coda running OK and data is good. Leave
           as is for now (entry number 307590 in Halog).
    
    04:55  Compton DAQ ROC0 not responding (for 15 mins). Coda reboot.
    
    06:20  Compton DAQ ROC6 not responding. Coda reboot several times
           with the same result being
    
    07:00  Asked MCC to stop beam, we move target to empty and request
           full radiological survey of Hall for restricted access to
           survey LHRS position and change RHRS angle.
    
    
    ==================================================================
    
    RUN LIST
    --------
    
    Compton
    -------
    runs 21635 thru 21640
    
    Happex  LHRS   RHRS   Time         Comments                 Good Pairs
    ------  -----  ----  -----------  ------------------------  ----------
    14717   26306  5811  23:20-00:17  production 105uA IHWP IN    33.2k
    14718   26307  5812  00:18-01:15  production 105uA IHWP IN    35.8k
    14719   26308  5813  01:16-01:51  production 105uA IHWP IN     3.2k
                                       (beam off for 25 mins)
    14720   26309  5814  01:55-02:53  production 105uA IHWP IN    37.1k
    14721   26310  5815  02:53-03:49  production 105uA IHWP IN    38.5k
    14722   26311  5816  03:59-04:55  production 105uA IHWP IN    29.0k
    14723   26312  5817  04:56-05:52  production 105uA IHWP IN    34.5k
    14724   26313  5818  05:54-06:50  production 105uA IHWP IN    37.1k 
    14725   26314  6819  06:50-07:02  production 105uA IHWP IN     4.6k
    ----------------------------------------------------------------------